# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 1.2 Setting
We recall the conventions and notations of [LLZ, §2]. Let $X$ be a smooth rigid analytic variety over $\mathrm{Spa}(K,\mathcal{O}_{K})$ and let $D\subset X$ be a strict normal crossing divisor with finitely many irreducible components $\{D_{j}\}_{j\in I}$ such that for every $J\subset I$
@@
In order to have a good crystalline theory we strengthen the assumptions of [LLZ] and we require the following :
@@ 2 Frobenius log-isocrystals
The Frobenius crystal $\bigl(\mathfrak{E},\Phi_{\mathfrak{E}}\bigr)$ defines a convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)$ on the log-convergent site of $\bigl(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\bigr)$ for the Zariski topology, relative to the thickening $\mathrm{Spec}(k)\to\mathcal{S}_{0}$ in the sense of [Shiho, Def. 2.1.5].
@@
with $(P,L)\to\mathcal{S}_{0}$ formally log-smooth endowed with a lift of Frobenius $\varphi_{P}$. Let $\mathcal{O}_{P}^{\rm DP}$ be the structure sheaf on $U$ of the $p$-adic DP envelope of $U\subset P$. Then, $\mathfrak{E}$ defines a locally free $\mathcal{O}_{P}^{\rm DP}$-module $\mathfrak{E}_{P}$ endowed with a logarithmic integrable connection that extends to an HPD stratification on $(U,\alpha_{U})$ relative to $(P,L)$ ; see the discussion in [ShihoI, §4.3]. The Frobenius structure provides a morphism $\Phi_{\mathfrak{E}_{P}}\colon\varphi_{P}^{\ast}(\mathfrak{E}_{P})\to\mathfrak{E}_{P}$, which is an isomorphism if we invert $p$. Arguing as in [Berthelot, §2.4.1] one deduces that the non-degenerate Frobenius log-crystal $\bigl(\mathfrak{E}_{P},\Phi_{\mathfrak{E}_{P}}\bigr)$ defines a module endowed with an integrable logarithmic connection, extending to a stratification, and an horizontal Frobenius $(\mathfrak{E}_{P}^{\rm conv},\Phi_{\mathfrak{E}_{P}^{\rm conv}})$ on the tubular neighborhood $U[^{\rm log}_{P}:=](U,\alpha_{U})[_{(P,L)}$. Recall from [Shiho, Def. 2.2.5] that $U[^{\rm log}_{P}$ is the rigid analytic fiber of the formal completion of $P$ along $U$. Since $U_{k}\subset U$ is a nilpotent thickening, the tubular neighborhoods $U_{k}[^{\rm log}_{P}$ and $U[^{\rm log}_{P}$ coincide. Hence, $(\mathfrak{E}_{P}^{\rm conv},\Phi_{\mathfrak{E}_{P}^{\rm conv}})$ defines a convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)|_{U}$ on the log-convergent site of $\bigl(U_{k},\alpha_{\mathcal{X}_{k}}|_{U_{k}}\bigr)$ relative to the thickening $\mathrm{Spec}(k)\to\mathcal{S}_{0}$ by [Shiho, Prop. 2.2.7]. Since convergent Frobenius isocrystals satisfy Zariski descent and for varying $U$’s the constructions above glue, we can glue and get the sought for convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)$. ∎
In [Shiho, Def. 2.2.12] the author defines complexes $\mathrm{R}\Gamma_{\mathrm{an}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}^{\rm conv}\bigr)$ computing the analytic cohomology of $\mathfrak{E}^{\rm conv}$. We sometimes omit $\alpha_{\mathcal{X}_{k}}$ from the notation for simplicity. The Frobenius defines $\sigma$-linear isomorphisms $\Phi$. We have a natural map, compatible with Frobenii, in the derived category
@@
is nilpotent. We recall that it is the morphism of $\mathcal{O}_{D_{J}}$ modules defined by composing $\nabla\colon\mathcal{E}\to\mathcal{E}\otimes_{\mathcal{O}_{X}}\Omega^{1}_{X}$ with the base change $\mathcal{E}\otimes_{\mathcal{O}_{X}}$ of the Poincaré residue map $\Omega^{1}_{X}\to\mathcal{O}_{D_{j}}$ along $D_{J}$ (see [OgusBook, Ex. 1.2.13 & Prop. 1.2.14]).
@@ 3 The syntomic complex
Note also that $\tau_{\rm crys}$ is an isomorphism in cohomology and, hence, the resulting syntomic cohomologies are the same when $K=K_{0}$ by [Shiho, Thm. 3.1.1], while in general the analytic cohomology computes the convergent cohomology by [Shiho, Thm. 2.3.9].
@@ 4.1 Chech covers
Given $\underline{\beta}:=(\beta_{1},\ldots,\beta_{n})\in B^{n}$ such that $\beta_{1}<\ldots<\beta_{n}$, we write $\mathcal{X}_{\underline{\beta}}:=\cap_{i=1}^{n}\mathcal{X}_{\beta_{i}}\subset\mathcal{X}$, $X_{\underline{\beta}}=\cap_{i=1}^{n}X_{\beta_{i}}\subset X$ and we let $\mathfrak{X}_{\underline{\beta}}\to\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ be the $(p,Z)$-adic formal scheme defined as the $(p,Z)$-adic completion of the log-envelope of the product $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ ; see [AI, Lemma 2.16]. It comes endowed with a log-structure $\widetilde{\alpha}_{\underline{\beta}}\colon\mathcal{M}_{\mathfrak{X}_{\underline{\beta}}}\to\mathcal{O}_{\mathfrak{X}_{\underline{\beta}}}$. Explicitly, given
@@
given by $X_{i,j}\mapsto v_{i,j}X_{i}$ and $Y_{i,k}\mapsto u_{i,k}Y_{k}$. The closed immersion $\mathcal{X}_{\underline{\beta}}\subset\prod_{i=1,\mathcal{S}}^{n}\mathcal{X}_{\beta_{i}}\subset\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ lifts to a closed immersion of $\iota_{\underline{\beta}}\colon\mathcal{X}_{\underline{\beta}}\to\mathfrak{X}_{\underline{\beta}}$. Furthermore, the product of the Fobenius lifts $\prod_{i=1}^{n}\varphi_{\beta_{i}}$ on $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ defines a lift $\varphi_{\underline{\beta}}$ of Frobenius to $\mathfrak{X}_{\underline{\beta}}$ by functoriality of the log-envelope. This also defines a morphism $\varphi_{\underline{\beta}}$ on the tubular neighborhood $\big]\mathcal{X}_{\underline{\beta},k}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}$ ; see [Shiho, Def. 2.2.5]. Notice that $\big]\mathcal{X}_{\underline{\beta},k}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}=\big]\overline{\mathcal{X}}_{\underline{\beta}}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}$ as $\mathcal{X}_{\underline{\beta},k}\subset\overline{\mathcal{X}}_{\underline{\beta}}$ is a closed immersion defined by a nilpotent ideal.
@@ 4.5 Log-rigid complexes
Consider $\underline{\beta}=(\beta_{1},\ldots,\beta_{n})\in B^{n}$ such that $\beta_{1}<\ldots<\beta_{n}$. Define $U_{\underline{\beta}}:=U\cap\mathcal{X}_{\underline{\beta},k}$ and $Y_{\underline{\beta}}:=Y\cap\mathcal{X}_{\underline{\beta},k}$. Restricting $(\mathfrak{E}^{\rm conv}_{\underline{\beta}},\nabla_{\underline{\beta}}^{\rm conv},\Phi_{\underline{\beta}}^{\rm conv})$ to strict neighborhoods of the tubular neighborhood of $U_{\underline{\beta}}$ in the tube of $Y_{\underline{\beta}}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$ we get a Frobenius overconvergent log-isocrystal $\mathfrak{E}_{\underline{\beta}}^{\dagger}$ relative to $U_{\underline{\beta}}\subset Y_{\underline{\beta}}$ and the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$. See [Shiho, Rmk. 2.2.19]. Proceeding as in the previous sections using open covers by quasi-Stein spaces as in Remark 4.1, we get a complex $\mathrm{R}\Gamma_{\mathrm{rig}}\bigl((U\subset Y,\alpha_{k})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ given by taking the total complex of the logarithmic de Rham complexes
@@
As the notation suggests, under the assumption that $U\subset\mathcal{X}_{k}^{\rm sm}$, the cohomology groups we obtain are independent of the choices made in §4.1. This is not clear in general. See [Shiho, Conj. 2.2.20].
@@
where $b$ is defined by exactifying the log-structure defined on each $\mathfrak{X}_{\beta_{i}}$ by the special fiber $\mathcal{X}_{\beta_{i},k}$ and $a$ is defined by exactifying the log-structure given on each $\mathfrak{X}_{\beta_{i}}$ by the divisor $\mathfrak{D}_{\beta_{i}}$. It follows from [Shiho, Claim p. 114] and the assumption that $U\subset\mathcal{X}_{k}^{\rm sm}$ that the morphism $b$ induces an isomorphism of log-rigid spaces $b_{V}\colon V_{\underline{\beta}}^{\log}\to V_{\underline{\beta}}$ on suitable strict neighborhoods $V_{\underline{\beta}}^{\log}$ and $V_{\underline{\beta}}$ of the tubular neighborhood of $U_{\underline{\beta}}$ in the tube of $\mathcal{X}_{\underline{\beta},k}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}^{\prime}$, resp. of $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$. Here, possibly enlarging the index set $B^{n}$, we can replace $V_{\underline{\beta}}$ with an open admissible cover by quasi-Stein spaces, see Remark 4.1, living over the rigid analytic fiber of $\mathfrak{S}$, with the log-structure pull-back of $\mathfrak{M}$.
@@ 5 Relation with Hyodo-Kato cohomology I
(2) Since each $\mathrm{H}^{j}_{\mathrm{HK}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0}^{0},\mathfrak{E}_{0}\bigr)$ is a finite dimensional $K_{0}$-vector space due to the properness assumption (see [BerthelotOgusBook, §7]), also each $\mathrm{H}^{j}_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}_{0}\bigr)$ is finite dimensional as $K_{0}$-vector space and, hence, as $\mathbb{Q}_{p}$-vector space.
@@
$\mathrm{ker}\left(H^{j}\left(y\right)\right)\overset{\sim}{\longrightarrow}\mathrm{ker}\left(H^{j}\left(z\right)\right)\text{ and }\mathrm{coker}\left(H^{j}\left(y\right)\right)\overset{\sim}{\longrightarrow}\mathrm{coker}\left(H^{j}\left(z\right)\right).$
@@ 6 Syntomic complexes with support conditions
The divisor $\mathcal{D}_{k}$ defines a crystal of invertible ideals $\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}(-\mathcal{D}_{k})\subset\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}$ ; see [EY, Def. 4.3 & Rmk, 4.5]. Given a Chech cover as in §4.1, for every $\beta\in B$ its value on $\mathfrak{X}_{\beta}$ is the ideal sheaf in $\mathcal{O}_{\mathfrak{X}_{\beta}}$ defined by the Cartier divisor $\mathfrak{D}_{\beta}\subset\mathfrak{X}_{\beta}$ of loc. cit. Let $\mathbf{1}(-D)\subset\mathbf{1}$ be the Frobenius subcrystal defined by $\mathcal{O}_{X}(-D)\subset\mathcal{O}_{X}$ and $\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}(-\mathcal{D}_{k})\subset\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}$. Given a non-degenerate Frobenius log-crystal $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\Phi_{\mathfrak{E}}\bigr)$ relative to $(X,\alpha_{X},\overline{\mathcal{X}},\alpha_{\overline{\mathcal{X}}})$ we write $\bigl(\mathcal{E}(-D),\mathrm{Fil}^{\bullet}\mathcal{E}(-D),\nabla,\mathfrak{E}(-\mathfrak{D}),\Phi_{\mathfrak{E}}(-\mathfrak{D})\bigr)$ for the tensor product with $\mathbf{1}(-D)$.
@@
Similarly we have the crystalline analytic syntomic complex $\mathbf{Syn}_{\rm crys}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)$ with support along $D$ with a morphism $\mathbf{Syn}_{\rm an}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}_{\rm crys}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)$ ; see Remark 3.2. One has also an analogue of Proposition 3.1.

### v2

@@ 1.2 Setting
We recall the conventions and notations of [LLZ23, §2]. Let $X$ be a smooth rigid analytic variety over $\mathrm{Spa}(K,\mathcal{O}_{K})$ and let $D\subset X$ be a strict normal crossing divisor with finitely many irreducible components $\{D_{j}\}_{j\in I}$ such that for every $J\subset I$
@@
In order to have a good crystalline theory we strengthen the assumptions of [LLZ23] and we require the following :
@@ 2 Frobenius log-isocrystals
The Frobenius crystal $\bigl(\mathfrak{E},\Phi_{\mathfrak{E}}\bigr)$ defines a convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)$ on the log-convergent site of $\bigl(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\bigr)$ for the Zariski topology, relative to the thickening $\mathrm{Spec}(k)\to\mathcal{S}_{0}$ in the sense of [SHI02, Def. 2.1.5].
@@
with $(P,L)\to\mathcal{S}_{0}$ formally log-smooth endowed with a lift of Frobenius $\varphi_{P}$. Let $\mathcal{O}_{P}^{\rm DP}$ be the structure sheaf on $U$ of the $p$-adic DP envelope of $U\subset P$. Then, $\mathfrak{E}$ defines a locally free $\mathcal{O}_{P}^{\rm DP}$-module $\mathfrak{E}_{P}$ endowed with a logarithmic integrable connection that extends to an HPD stratification on $(U,\alpha_{U})$ relative to $(P,L)$ ; see the discussion in [SHI00, §4.3]. The Frobenius structure provides a morphism $\Phi_{\mathfrak{E}_{P}}\colon\varphi_{P}^{\ast}(\mathfrak{E}_{P})\to\mathfrak{E}_{P}$, which is an isomorphism if we invert $p$. Arguing as in [BER96, §2.4.1] one deduces that the non-degenerate Frobenius log-crystal $\bigl(\mathfrak{E}_{P},\Phi_{\mathfrak{E}_{P}}\bigr)$ defines a module endowed with an integrable logarithmic connection, extending to a stratification, and an horizontal Frobenius $(\mathfrak{E}_{P}^{\rm conv},\Phi_{\mathfrak{E}_{P}^{\rm conv}})$ on the tubular neighborhood $U[^{\rm log}_{P}:=](U,\alpha_{U})[_{(P,L)}$. Recall from [SHI02, Def. 2.2.5] that $U[^{\rm log}_{P}$ is the rigid analytic fiber of the formal completion of $P$ along $U$. Since $U_{k}\subset U$ is a nilpotent thickening, the tubular neighborhoods $U_{k}[^{\rm log}_{P}$ and $U[^{\rm log}_{P}$ coincide. Hence, $(\mathfrak{E}_{P}^{\rm conv},\Phi_{\mathfrak{E}_{P}^{\rm conv}})$ defines a convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)|_{U}$ on the log-convergent site of $\bigl(U_{k},\alpha_{\mathcal{X}_{k}}|_{U_{k}}\bigr)$ relative to the thickening $\mathrm{Spec}(k)\to\mathcal{S}_{0}$ by [SHI02, Prop. 2.2.7]. Since convergent Frobenius isocrystals satisfy Zariski descent and for varying $U$’s the constructions above glue, we can glue and get the sought for convergent Frobenius isocrystal $\bigl(\mathfrak{E}^{\rm conv},\Phi_{\mathfrak{E}^{\rm conv}}\bigr)$. ∎
In [SHI02, Def. 2.2.12] the author defines complexes $\mathrm{R}\Gamma_{\mathrm{an}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}^{\rm conv}\bigr)$ computing the analytic cohomology of $\mathfrak{E}^{\rm conv}$. We sometimes omit $\alpha_{\mathcal{X}_{k}}$ from the notation for simplicity. The Frobenius defines $\sigma$-linear isomorphisms $\Phi$. We have a natural map, compatible with Frobenii, in the derived category
@@
is nilpotent. We recall that it is the morphism of $\mathcal{O}_{D_{J}}$ modules defined by composing $\nabla\colon\mathcal{E}\to\mathcal{E}\otimes_{\mathcal{O}_{X}}\Omega^{1}_{X}$ with the base change $\mathcal{E}\otimes_{\mathcal{O}_{X}}$ of the Poincaré residue map $\Omega^{1}_{X}\to\mathcal{O}_{D_{j}}$ along $D_{J}$ (see [OGU18, Ex. 1.2.13 & Prop. 1.2.14]).
@@ 3 The syntomic complex
Note also that $\tau_{\rm crys}$ is an isomorphism in cohomology and, hence, the resulting syntomic cohomologies are the same when $K=K_{0}$ by [SHI02, Thm. 3.1.1], while in general the analytic cohomology computes the convergent cohomology by [SHI02, Thm. 2.3.9].
@@ 4.1 Chech covers
Given $\underline{\beta}:=(\beta_{1},\ldots,\beta_{n})\in B^{n}$ such that $\beta_{1}<\ldots<\beta_{n}$, we write $\mathcal{X}_{\underline{\beta}}:=\cap_{i=1}^{n}\mathcal{X}_{\beta_{i}}\subset\mathcal{X}$, $X_{\underline{\beta}}=\cap_{i=1}^{n}X_{\beta_{i}}\subset X$ and we let $\mathfrak{X}_{\underline{\beta}}\to\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ be the $(p,Z)$-adic formal scheme defined as the $(p,Z)$-adic completion of the log-envelope of the product $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ ; see [AI12, Lemma 2.16]. It comes endowed with a log-structure $\widetilde{\alpha}_{\underline{\beta}}\colon\mathcal{M}_{\mathfrak{X}_{\underline{\beta}}}\to\mathcal{O}_{\mathfrak{X}_{\underline{\beta}}}$. Explicitly, given
@@
given by $X_{i,j}\mapsto v_{i,j}X_{i}$ and $Y_{i,k}\mapsto u_{i,k}Y_{k}$. The closed immersion $\mathcal{X}_{\underline{\beta}}\subset\prod_{i=1,\mathcal{S}}^{n}\mathcal{X}_{\beta_{i}}\subset\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ lifts to a closed immersion of $\iota_{\underline{\beta}}\colon\mathcal{X}_{\underline{\beta}}\to\mathfrak{X}_{\underline{\beta}}$. Furthermore, the product of the Fobenius lifts $\prod_{i=1}^{n}\varphi_{\beta_{i}}$ on $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$ defines a lift $\varphi_{\underline{\beta}}$ of Frobenius to $\mathfrak{X}_{\underline{\beta}}$ by functoriality of the log-envelope. This also defines a morphism $\varphi_{\underline{\beta}}$ on the tubular neighborhood $\big]\mathcal{X}_{\underline{\beta},k}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}$ ; see [SHI02, Def. 2.2.5]. Notice that $\big]\mathcal{X}_{\underline{\beta},k}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}=\big]\overline{\mathcal{X}}_{\underline{\beta}}\big[_{\mathfrak{X}_{\underline{\beta}}}^{\log}$ as $\mathcal{X}_{\underline{\beta},k}\subset\overline{\mathcal{X}}_{\underline{\beta}}$ is a closed immersion defined by a nilpotent ideal.
@@ 4.5 Log-rigid complexes
Consider $\underline{\beta}=(\beta_{1},\ldots,\beta_{n})\in B^{n}$ such that $\beta_{1}<\ldots<\beta_{n}$. Define $U_{\underline{\beta}}:=U\cap\mathcal{X}_{\underline{\beta},k}$ and $Y_{\underline{\beta}}:=Y\cap\mathcal{X}_{\underline{\beta},k}$. Restricting $(\mathfrak{E}^{\rm conv}_{\underline{\beta}},\nabla_{\underline{\beta}}^{\rm conv},\Phi_{\underline{\beta}}^{\rm conv})$ to strict neighborhoods of the tubular neighborhood of $U_{\underline{\beta}}$ in the tube of $Y_{\underline{\beta}}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$ we get a Frobenius overconvergent log-isocrystal $\mathfrak{E}_{\underline{\beta}}^{\dagger}$ relative to $U_{\underline{\beta}}\subset Y_{\underline{\beta}}$ and the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$. See [SHI02, Rmk. 2.2.19]. Proceeding as in the previous sections using open covers by quasi-Stein spaces as in Remark 4.1, we get a complex $\mathrm{R}\Gamma_{\mathrm{rig}}\bigl((U\subset Y,\alpha_{k})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ given by taking the total complex of the logarithmic de Rham complexes
@@
As the notation suggests, under the assumption that $U\subset\mathcal{X}_{k}^{\rm sm}$, the cohomology groups we obtain are independent of the choices made in §4.1. This is not clear in general. See [SHI02, Conj. 2.2.20].
@@
where $b$ is defined by exactifying the log-structure defined on each $\mathfrak{X}_{\beta_{i}}$ by the special fiber $\mathcal{X}_{\beta_{i},k}$ and $a$ is defined by exactifying the log-structure given on each $\mathfrak{X}_{\beta_{i}}$ by the divisor $\mathfrak{D}_{\beta_{i}}$. It follows from [SHI02, Claim p. 114] and the assumption that $U\subset\mathcal{X}_{k}^{\rm sm}$ that the morphism $b$ induces an isomorphism of log-rigid spaces $b_{V}\colon V_{\underline{\beta}}^{\log}\to V_{\underline{\beta}}$ on suitable strict neighborhoods $V_{\underline{\beta}}^{\log}$ and $V_{\underline{\beta}}$ of the tubular neighborhood of $U_{\underline{\beta}}$ in the tube of $\mathcal{X}_{\underline{\beta},k}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}^{\prime}$, resp. of $\prod_{i=1,\mathfrak{S}}^{n}\mathfrak{X}_{\beta_{i}}$. Here, possibly enlarging the index set $B^{n}$, we can replace $V_{\underline{\beta}}$ with an open admissible cover by quasi-Stein spaces, see Remark 4.1, living over the rigid analytic fiber of $\mathfrak{S}$, with the log-structure pull-back of $\mathfrak{M}$.
@@ 5 Relation with Hyodo-Kato cohomology I
(2) Since each $\mathrm{H}^{j}_{\mathrm{HK}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0}^{0},\mathfrak{E}_{0}\bigr)$ is a finite dimensional $K_{0}$-vector space due to the properness assumption (see [BO78, §7]), also each $\mathrm{H}^{j}_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}_{0}\bigr)$ is finite dimensional as $K_{0}$-vector space and, hence, as $\mathbb{Q}_{p}$-vector space.
@@
$\mathrm{ker}\left(H^{j}\left(y\right)\right)\overset{\sim}{\longrightarrow}\mathrm{ker}\left(H^{j}\left(z\right)\right)\text{ and }\mathrm{coker}\left(H^{j}\left(y\right)\right)\overset{\sim}{\longrightarrow}\mathrm{coker}\left(H^{j}\left(z\right)\right)\text{.}$
@@ 6 Syntomic complexes with support conditions
The divisor $\mathcal{D}_{k}$ defines a crystal of invertible ideals $\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}(-\mathcal{D}_{k})\subset\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}$ ; see [EY20, Def. 4.3 & Rmk, 4.5]. Given a Chech cover as in §4.1, for every $\beta\in B$ its value on $\mathfrak{X}_{\beta}$ is the ideal sheaf in $\mathcal{O}_{\mathfrak{X}_{\beta}}$ defined by the Cartier divisor $\mathfrak{D}_{\beta}\subset\mathfrak{X}_{\beta}$ of loc. cit. Let $\mathbf{1}(-D)\subset\mathbf{1}$ be the Frobenius subcrystal defined by $\mathcal{O}_{X}(-D)\subset\mathcal{O}_{X}$ and $\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}(-\mathcal{D}_{k})\subset\mathcal{O}_{\mathcal{X}_{k}}^{\rm crys}$. Given a non-degenerate Frobenius log-crystal $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\Phi_{\mathfrak{E}}\bigr)$ relative to $(X,\alpha_{X},\overline{\mathcal{X}},\alpha_{\overline{\mathcal{X}}})$ we write $\bigl(\mathcal{E}(-D),\mathrm{Fil}^{\bullet}\mathcal{E}(-D),\nabla,\mathfrak{E}(-\mathfrak{D}),\Phi_{\mathfrak{E}}(-\mathfrak{D})\bigr)$ for the tensor product with $\mathbf{1}(-D)$.
@@
Similarly we have the crystalline analytic syntomic complex $\mathbf{Syn}_{\rm crys}^{c-\infty}(\mathcal{E},\mathfrak{E},r)$ with support along $D$ with a morphism $\mathbf{Syn}_{\rm an}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}_{\rm crys}^{c-\infty}(\mathcal{E},\mathfrak{E},r)$ ; see Remark 3.2. One has also an analogue of Proposition 3.1.